A particular brand of beef jerky contains 0.0552 % sodium nitrite (NaNO2) by mass and is sold in an 8 oz (1 oz = 28.35 g) bag. What mass of sodium does the sodium nitrite contribute to the bag of beef jerky?


1
Expert's answer
2021-09-06T02:37:57-0400

Total mass=(8*28.35g)=226.8g

Mass of NaNO2= (0.0552*226.8/100)=0.1251936g

Moles of NaNO2=(0.1251936/68.9953)=0.00181452moles

NaNO2(aq)=Na-(aq) + NO2+(aq)

Moles ratio=1:1:1

Moles of Na-(aq)=0.00181452moles

Mass of Na=(0.00181452*22.989769)

=0.04172 g of Na